Okay so today in math my teacher taught us on evaluating expressions and like she wants us to show summ kind of work like:

6 to the power of 5 so you go
6*6*6*6*6
Awnswer is 7796
then she wants us to do negatives and i dont really get how to do those??
Could you help?? Please thanks tesha

Answer by checkley77(12569)

First we handle the range of exponents thus:
5^3=5*5*5=125
5^2=5*5=25
5^1=5
5^0=1 *by definition)
5^(-1)=1/5
5^(-2)=1/5*5=1/25
5^(-3)=1/5*5*5=1/125
Also:
1/5^(-2)=1/(1/25)=1*25/1=25 or 5^2
5^5/5^(-2)=5^5*5^2=5^7
Hope this helps.